OpenStudy (anonymous):

Write the slope-intercept equation for the line with slope 1/2 and y-intercept (0, -5) and graph the line. Please show all of your work.

OpenStudy (anonymous):

y=1/2x - 5 5=1/2x x=10 y intercept -5 x intercept 10